OpenStudy (anonymous):

Solve using the square root method. x2 - 16 = 0 A. {-8, 8} B. {-4, 4} C. {8} D. {4} Solve using the square root method. x2 - 16 = 0 A. {-8, 8} B. {-4, 4} C. {8} D. {4} @Mathematics

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Do you understand why it is (a)??

OpenStudy (anonymous):

Solution to x^2-16=0 is to be computed as follows: x^2 = 16 So, x = square root of (16). Square root of 16 can have one of two values. Positive 4 and negative 4. Squaring both will result in 16. So, the answer is B.

OpenStudy (anonymous):

just square both sides of equation so.. sqrt of x^2 = x, sqrt of 16 is + or -4 so x = + or -4, B